# Gravitational Positivity Bounds on Higgs-Portal Light Dark Matter

### v2

@@ 4 Phenomenological Implications for Light Dark Matter
For the parameter set given in Eq. (34), and with $\Lambda\lesssim 10^{16}$ GeV, the possible additional contribution on the RHS of Eq. (17), which could potentially lead to a positivity violation, is of order $\mathcal{O}\bigl(1/(\Lambda^{2}_{\mathrm{QG}}M^{2}_{\text{pl}})\bigr)=\mathcal{O}(\alpha^{\prime}/M^{2}_{\text{pl}})$ on the RHS of Eq. (17) can be safely neglected. Here, we assume a typical value of $\Lambda_{\mathrm{QG}}=\alpha^{\prime-1/2}\sim 10^{16}$ GeV, representing the scale of Reggeization [35, 56, 12, 48].
Finally, we comment on the gravitational contribution $B_{\mathrm{grav}}$ associated with the four-point self-interaction of DM. In this case, the one-loop contribution is absent, and the leading contribution arises at the two-loop level [48]. As a result, $B_{\mathrm{grav}}$ itself, which provides a lower bound, receives an additional contribution scaling as $\mathcal{O}\!\left(\lambda_{\phi}^{2}/(m^{2}_{\phi}\overline{M}^{2}_{\mathrm{pl}})\right)$ [48]. For sub-GeV DM with a GUT-scale (or lower) cutoff, however, this contribution can be safely neglected in comparison with the non-gravitational contribution from the self-coupling, which scales as $\sim\lambda_{\phi}^{2}/\Lambda^{4}$ in $B_{\mathrm{non\text{-}grav}}$. For DM masses lighter than the sub-GeV scale, the positivity bounds become more stringent, but this does not alter our conclusion, as our bound is conservative. For heavier DM masses, the two-loop gravitational contribution is further suppressed and can therefore be neglected.
@@ 5 Summary and Discussion
We have considered a Higgs-portal real scalar DM model with a $Z_{2}$-odd parity and renormalizable interactions. In this framework, we applied gravitational positivity bounds to the forward DM scattering process $\phi\phi\to\phi\phi$. For light DM with $m_{\phi}<m_{h}$ and in the absence of a DM self-coupling, new physics beyond the Higgs-portal scenario is required to appear below an energy scale of $10^{10}$ GeV.
When a DM self-coupling is present, a large hierarchy between the self-coupling $\lambda_{\phi}$ and the Higgs-portal coupling $\lambda_{h\phi}$ raises the upper bound on the new physics scale $\Lambda$. For example, with $\lambda_{\phi}/\lambda_{h\phi}=10^{12}$, the gravitational positivity bound allows a $\Lambda$ at the GUT scale or at a typical string scale. In such a scenario, the dark freeze-out mechanism can still account for the observed relic abundance of DM in the Universe. We find that parameter values of $\lambda_{\phi}\sim\mathcal{O}(1)$, $\lambda_{h\phi}\sim 10^{-12}$, and sub-GeV DM are consistent with achieving a GUT-scale or string-scale $\Lambda$ within the Higgs-portal light DM framework.
@@ Acknowledgements
To obtain the relevant part of Eq. (7), we multiply $(s-2m^{2}_{\phi}+t/2)^{2}$ to the above Eq. (37). Here,

### v3

@@ 3.2 Gravitational Positivity Bounds with $\phi\phi\to\phi\phi$
Let us comment on the possible additional contribution on the RHS of Eq. (17), which could potentially lead to a violation of the positivity bound. When setting the GUT scale as the cutoff scale, $\Lambda=\Lambda_{\text{GUT}}$, the contribution of order $\mathcal{O}\!\left(1/(\Lambda_{\mathrm{QG}}^{2}\overline{M}^{2}_{\text{pl}})\right)=\mathcal{O}(\alpha^{\prime}/\overline{M}^{2}_{\text{pl}})$ on the RHS of Eq. (17) can be safely neglected. Here we assume a typical value of $\Lambda_{\mathrm{QG}}=\alpha^{\prime-1/2}\sim 10^{16}\,\mathrm{GeV}$, which represents the scale of Reggeization [15, 14, 16, 17]. In this case, the applicability condition
$\min(\Lambda,\Lambda_{*})<\sqrt{\overline{M}_{\text{pl}}\,\Lambda_{\mathrm{QG}}}$
is satisfied, where $\Lambda_{*}=B_{\text{grav}}^{-1/4}$ [17].
4 Phenomenological Implications for Heavy Dark Matter
Based on the results presented in Sec. 3.2, we discuss the phenomenological implications of the gravitational positivity bounds for the Higgs-portal real scalar DM model in the heavy-mass regime. If the new physics scale $\Lambda$ is identified with the GUT scale or a typical string scale, $\Lambda\sim 10^{16}\,\mathrm{GeV}$, the bound implies a DM mass of order $10^{10}$$10^{11}\,\mathrm{GeV}$ (or above).
The Higgs-portal coupling is constrained from the perspective of the DM relic abundance. There may exist additional scattering processes that lead to further gravitational positivity bounds, for example by constraining the couplings, which we leave such analyses for future work.
Since the DM mass is extremely large, the conventional weakly interacting massive particle (WIMP) scenario cannot reproduce the observed relic abundance due to the perturbative unitarity bound on the Higgs-portal coupling.
We therefore consider the feebly interacting massive particle (FIMP) scenario, in which DM never thermalizes with the SM plasma, following Ref. [51]. Here we focus on DM production after reheating. DM production during the reheating era depends on the inflationary dynamics and sets the initial condition for subsequent production.
$T_{\mathrm{reh}}>m_{\phi}$ and the DM is heavy, gravitational interactions can also contribute to DM production in addition to the Higgs-portal interaction. We therefore include this contribution in our analysis.
Gravitational interactions are described by
$\mathcal{L}_{\text{grav}}=\frac{1}{\overline{M}_{\text{pl}}}h^{\mu\nu}\left(T^{\mathrm{SM}}_{\mu\nu}+T^{h}_{\mu\nu}+T^{\phi}_{\mu\nu}\right),$ (34)
where $h^{\mu\nu}$ denotes the graviton field. The energy–momentum tensor for a real scalar field $S$ is given by
$T^{S}_{\mu\nu}=\partial_{\mu}S\,\partial_{\nu}S-\frac{1}{2}\eta_{\mu\nu}\eta^{\rho\sigma}\partial_{\rho}S\,\partial_{\sigma}S-\frac{1}{2}\eta_{\mu\nu}m_{S}^{2}S^{2},\qquad(S=h,\,\phi).$ (35)
Based on the interaction in Eq. (34), the scalar DM particle $\phi$ can be produced from the SM plasma. The scattering amplitude for the process $h+h\to\phi+\phi$ via graviton exchange is given by [51]
$\mathcal{M}^{G}_{h+h\to\phi+\phi}=-\frac{1}{\overline{M}_{\text{pl}}^{2}}\frac{(t-m_{\phi}^{2})(s+t-m_{\phi}^{2})}{s},$ (36)
where $s$ and $t$ are the Mandelstam variables. Including the Higgs-portal interaction, the total squared amplitude for $h+h\to\phi+\phi$ becomes
$\left|\mathcal{M}^{\mathrm{total}}_{h+h\to\phi+\phi}\right|^{2}=\left(\lambda_{h\phi}+\frac{(t-m_{\phi}^{2})(s+t-m_{\phi}^{2})}{s\,\overline{M}_{\text{pl}}^{2}}\right)^{2}.$ (37)
Similarly, the contributions from other SM particles through graviton exchange are given by [51]
$\left|\mathcal{M}^{G}_{f+f\to\phi+\phi}\right|^{2}$ $=-\frac{1}{2\overline{M}_{\text{pl}}^{4}s^{2}}\left(s+2t-2m_{\phi}^{2}\right)^{2}\left[\left(t-m_{\phi}^{2}\right)^{2}+st\right],$ (38)
$\left|\mathcal{M}^{G}_{V+V\to\phi+\phi}\right|^{2}$ $=\frac{2}{\overline{M}_{\text{pl}}^{4}s^{2}}\left(m_{\phi}^{4}-2m_{\phi}^{2}t+t(s+t)\right)^{2},$ (39)
where $f$ and $V$ denote SM fermions and SM vector bosons, respectively.
When DM is decoupled from the SM plasma, its number density $n_{\phi}$ evolves according to the Boltzmann equation with a production reaction rate [52]
$\dot{n}_{\phi}+3Hn_{\phi}=R(T),$ (40)
where $R(T)$ denotes the reaction rate for thermal scattering processes.
For the thermal production process $i_{1}(p_{1})+i_{2}(p_{2})\to\phi(p_{3})+\phi(p_{4})$, the reaction rate can be written in terms of the squared amplitude $\overline{|\mathcal{M}_{i_{1}+i_{2}\to\phi+\phi}|^{2}}$ as [52, 53]
$R(T)=\frac{T}{2^{11}\pi^{6}}\int_{4m_{\phi}^{2}}^{\infty}\!ds\int d\Omega\,K_{1}\!\left(\frac{\sqrt{s}}{T}\right)\sqrt{s-4m_{\phi}^{2}}\,\overline{|\mathcal{M}_{i_{1}+i_{2}\to\phi+\phi}|^{2}},$ (41)
where $i_{1,2}$ collectively denote the SM radiation. The solid angle element is defined as $d\Omega\equiv 2\pi\,d\cos\theta_{13}$, with $\theta_{13}$ being the angle between the momenta $\mathbf{p}_{1}$ and $\mathbf{p}_{3}$ in the center-of-mass frame, and $K_{1}(z)$ is the modified Bessel function of the second kind. The overline indicates that appropriate symmetry factors for the initial and final states are included.
After reheating, using $T\propto a^{-1}$ (and hence $\dot{T}=-HT$) together with
$H=\sqrt{\frac{g_{\mathrm{reh}}\pi^{2}}{90}}\frac{T^{2}}{\overline{M}_{\text{pl}}},$ (42)
where $g_{\mathrm{reh}}=106.75$ denotes the effective number of relativistic degrees of freedom during reheating, the Boltzmann equation (40) can be rewritten in terms of the DM yield $Y\equiv n_{\phi}/T^{3}$ as
$\frac{dY}{dT}=-\frac{1}{HT^{4}}R(T)=-\sqrt{\frac{90}{\pi^{2}g_{\mathrm{reh}}}}\frac{\overline{M}_{\text{pl}}}{T^{6}}R(T).$ (43)
The total reaction rate can be expressed as
$R_{\rm{total}}(T)=4R_{h}(T)+45R_{f}(T)+12R_{V}(T),$ (44)
where $R_{h}$, $R_{f}$, and $R_{V}$ are the reaction rates defined in Eq. (41), corresponding to the matrix elements given in Eqs. (37), (38), and (39), respectively. The numerical coefficients account for the relevant SM degrees of freedom. Here we consider the Higgs-unbroken phase.
From Ref. [51], the combined contribution from SM fermions and vector bosons is given by
$45R_{f}(T)+12R_{V}(T)=\frac{69m_{\phi}^{7}T}{512\pi^{9/2}M^{4}_{\text{pl}}}G_{1,3}^{3,0}\left.\left(\begin{array}[]{c}0\\ -\frac{7}{2},-\frac{1}{2},\frac{1}{2}\end{array}\right|\frac{m_{\phi}^{2}}{T^{2}}\right),$ (47)
where $G_{p,q}^{m,n}\left.\left(\begin{array}[]{c}a_{1},\ldots,a_{p}\\ b_{1},\ldots,b_{q}\end{array}\right|z\right)$ denotes the Meijer $G$-function.
We obtain the Higgs contribution
$4R_{h}(T)$ $=\frac{\lambda_{h\phi}^{2}\,m_{\phi}^{2}T^{2}}{128\,\pi^{5}}K_{1}^{2}\!\left(\frac{m_{\phi}}{T}\right)$
$\quad+\frac{m_{\phi}^{2}T}{1920\,\pi^{5}\overline{M}_{\text{pl}}^{4}}\Biggl[4\sqrt{\pi}\,m_{\phi}^{5}\,G_{1,3}^{3,0}\!\left(\left.\begin{array}[]{c}-2\\ -\frac{7}{2},\,-\frac{1}{2},\,\frac{1}{2}\end{array}\right|\frac{m_{\phi}^{2}}{T^{2}}\right)$ (50)
+ $\qquad\qquad+\,2\sqrt{\pi}\,m_{\phi}^{5}\,G_{1,3}^{3,0}\!\left(\left.\begin{array}[]{c}-1\\ -\frac{5}{2},\,-\frac{1}{2},\,\frac{1}{2}\end{array}\right|\frac{m_{\phi}^{2}}{T^{2}}\right)+\,3m_{\phi}^{4}T\,K_{1}^{2}\!\left(\frac{m_{\phi}}{T}\right)\Biggr]$ (53)
$\quad+\frac{\lambda_{h\phi}\,m_{\phi}^{2}T}{1920\,\pi^{5}\overline{M}_{\text{pl}}^{4}}\Biggl[-10\sqrt{\pi}\,\overline{M}_{\text{pl}}^{2}m_{\phi}^{3}\,G_{1,3}^{3,0}\!\left(\left.\begin{array}[]{c}-1\\ -\frac{5}{2},\,-\frac{1}{2},\,\frac{1}{2}\end{array}\right|\frac{m_{\phi}^{2}}{T^{2}}\right)$ (56)
$\qquad\qquad-10\overline{M}_{\text{pl}}^{2}m_{\phi}^{2}T\,K_{1}^{2}\!\left(\frac{m_{\phi}}{T}\right)\Biggr].$ (57)
The first term in Eq. (57) originates from the Higgs-portal interaction contribution to $hh\to\phi\phi$. The second square-bracketed term arises from purely gravitational interactions, and the third square-bracketed term corresponds to their interference.
Combining Eqs. (47) and (57), the total reaction rate in Eq. (44) can be written as
$R_{\text{tot}}(T)$ $=R_{\text{portal}}(T)+R_{\text{grav}}(T)+R_{\text{int}}(T),$ (58)
$R_{\text{portal}}(T)$ $=\frac{\lambda_{h\phi}^{2}\,m_{\phi}^{2}T^{2}}{128\,\pi^{5}}\,K_{1}^{2}\!\left(\frac{m_{\phi}}{T}\right),$ (59)
$R_{\text{grav}}(T)$ $=\frac{m_{\phi}^{2}T}{1920\,\pi^{5}\overline{M}_{\text{pl}}^{4}}\Biggl[4\sqrt{\pi}\,m_{\phi}^{5}\,G_{1,3}^{3,0}\!\left(\left.\begin{array}[]{c}-2\\ -\frac{7}{2},\,-\frac{1}{2},\,\frac{1}{2}\end{array}\right|\frac{m_{\phi}^{2}}{T^{2}}\right)$ (62)
$\qquad+2\sqrt{\pi}\,m_{\phi}^{5}\,G_{1,3}^{3,0}\!\left(\left.\begin{array}[]{c}-1\\ -\frac{5}{2},\,-\frac{1}{2},\,\frac{1}{2}\end{array}\right|\frac{m_{\phi}^{2}}{T^{2}}\right)+3m_{\phi}^{4}T\,K_{1}^{2}\!\left(\frac{m_{\phi}}{T}\right)\Biggr]$ (65)
$\quad+\frac{69\,m_{\phi}^{7}T}{512\,\pi^{9/2}\overline{M}_{\text{pl}}^{4}}G_{1,3}^{3,0}\!\left(\left.\begin{array}[]{c}0\\ -\frac{7}{2},\,-\frac{1}{2},\,\frac{1}{2}\end{array}\right|\frac{m_{\phi}^{2}}{T^{2}}\right),$ (68)
$R_{\text{int}}(T)$ $=\frac{\lambda_{h\phi}\,m_{\phi}^{2}T}{1920\,\pi^{5}\overline{M}_{\text{pl}}^{4}}\Biggl[-10\sqrt{\pi}\,\overline{M}_{\text{pl}}^{2}m_{\phi}^{3}\,G_{1,3}^{3,0}\!\left(\left.\begin{array}[]{c}-1\\ -\frac{5}{2},\,-\frac{1}{2},\,\frac{1}{2}\end{array}\right|\frac{m_{\phi}^{2}}{T^{2}}\right)$ (71)
$\qquad-10\overline{M}_{\text{pl}}^{2}m_{\phi}^{2}T\,K_{1}^{2}\!\left(\frac{m_{\phi}}{T}\right)\Biggr].$ (72)
Here $R_{\text{portal}}(T)$ arises from the Higgs-portal interaction, $R_{\text{grav}}(T)$ collects purely gravitational contributions from the Higgs sector and other SM particles, and $R_{\text{int}}(T)$ represents the interference between the Higgs-portal and gravitational amplitudes.
Integrating Eq. (43) over temperature from $T_{\mathrm{reh}}$ down to $T_{*}$, with the hierarchy $T_{*}\ll m_{X}\ll T_{\mathrm{reh}}$, we obtain a late-time DM yield that no longer depends on the choice of $T_{*}$. The resulting asymptotic abundance can be written as
$Y(T_{*})\simeq Y(T_{\mathrm{reh}})+\lambda_{h\phi}^{2}\frac{9\sqrt{10}\,\overline{M}_{\text{pl}}}{4096\,\pi^{4}\,g_{\text{reh}}^{1/2}\,m_{\phi}}+\frac{73\sqrt{10}\,T_{\text{reh}}^{3}}{80\,\pi^{6}\,g_{\text{reh}}^{1/2}\,\overline{M}_{\text{pl}}^{3}}+\lambda_{h\phi}\frac{\sqrt{10}\,T_{\text{reh}}}{16\,\pi^{6}\,g_{\text{reh}}^{1/2}\,\overline{M}_{\text{pl}}}.$ (73)
In this expression, $g_{\mathrm{reh}}$ has been taken to be constant throughout the integration. The value of the yield at the reheating temperature, $Y(T_{\mathrm{reh}})$, is sensitive to the initial conditions and to the detailed dynamics during the reheating era and we put $Y(T_{\mathrm{reh}})=0$ in our analysis.
Using the asymptotic value of the yield in Eq. (73), the present-day DM relic abundance can be evaluated as
$\Omega h^{2}=1.6\times 10^{8}\left(\frac{m_{\phi}}{1\,\mathrm{GeV}}\right)\left(\frac{g_{0}}{g_{\mathrm{reh}}}\right)\Biggl[\lambda_{h\phi}^{2}\frac{9\sqrt{10}\,\overline{M}_{\text{pl}}}{4096\,\pi^{4}\,g_{\text{reh}}^{1/2}\,m_{\phi}}\\ +\frac{73\sqrt{10}\,T_{\mathrm{reh}}^{3}}{80\,\pi^{6}\,g_{\text{reh}}^{1/2}\,\overline{M}_{\text{pl}}^{3}}+\lambda_{h\phi}\frac{\sqrt{10}\,T_{\mathrm{reh}}}{16\,\pi^{6}\,g_{\text{reh}}^{1/2}\,\overline{M}_{\text{pl}}}\Biggr].$ (74)
Here $g_{0}=3.91$ denotes the effective number of relativistic degrees of freedom at the present epoch. The first term in Eq. (74) corresponds to IR-dominated freeze-in via the Higgs-portal interaction and is controlled by temperatures of order $T\sim m_{\phi}$. The second term represents UV-dominated freeze-in due to graviton exchange and is governed by the highest temperature, leading to a characteristic $T_{\mathrm{reh}}^{3}$ dependence. The third term arises from the interference between the Higgs-portal and gravitational amplitudes, yielding a linear dependence on $T_{\mathrm{reh}}$.
Figure 9: Parameter space in $(m_{\phi},T_{\mathrm{reh}})$ for the DM abundance. The blue shaded region indicates $\Omega h^{2}>0.12$. The region with $T_{\mathrm{reh}}>m_{\phi}$ in the figure (namely, the region above the gray dashed line) is valid under the current approximation (73).
In Figure 9, we show the parameter space in $(m_{\phi},T_{\mathrm{reh}})$ under the condition that the observed DM abundance is saturated, $\Omega h^{2}=0.12$. The green dotted line in Figure 9 represents the contribution from gravitational interactions (corresponding to the second term in Eq. (74)), whereas the black dashed and solid lines show the total DM production, including the Higgs-portal interaction (from the first and third terms in Eq. (74)), for $\lambda_{h\phi}=3.0\times 10^{-11}$ (black dashed) and $\lambda_{h\phi}=3.5\times 10^{-11}$ (black solid), respectively. The blue shaded region indicates DM overclosure, namely $\Omega h^{2}>0.12$. We note that only the region with $T_{\mathrm{reh}}\gg m_{\phi}$ in the figure (i.e., above the gray dashed line) is consistent with the assumption used in Eq. (73).
Therefore, Higgs-portal scalar DM over a broad mass range can account for the entire observed DM abundance by varying the reheating temperature. However, in order to avoid DM overproduction, the Higgs-portal coupling must be tiny, $\lambda_{h\phi}\lesssim 3.5\times 10^{-11}$. We also note that for $\lambda_{h\phi}=3.6\times 10^{-11}$, the first term in Eq. (74) yields $\Omega h^{2}\simeq 0.13$, which exceeds the observed value. When only the Higgs-portal interaction is taken into account, the DM abundance is independent of both the DM mass and the reheating temperature.
From Sec. 3.2, DM masses of order $10^{10}$$10^{11}\,\mathrm{GeV}$ (or larger) are required to realize a GUT-scale cutoff. Consequently, eheating temperatures, $T_{\mathrm{reh}}\lesssim 10^{14}\,\mathrm{GeV}$, are required in Figure 9.
In this analysis, we set the initial DM abundance (the amount of DM produced at the end of the reheating epoch) to zero, i.e., $Y(T_{\mathrm{reh}})=0$ in Eq. (73). This initial condition depends on the inflaton dynamics during reheating. However, since a nonzero initial abundance only increases the final DM relic abundance, the bound $T_{\mathrm{reh}}\lesssim 10^{14}\,\mathrm{GeV}$ remains valid, as a higher reheating temperature enhances DM production from graviton-mediated interactions.
@@ 5 Summary and Discussion
We have considered a Higgs-portal real scalar DM model with a $Z_{2}$-odd parity and renormalizable interactions. In this framework, we applied gravitational positivity bounds to the forward DM scattering process $\phi\phi\to\phi\phi$.
For light DM with $m_{\phi}<m_{h}$ and in the absence of a DM self-coupling, new physics beyond the Higgs-portal scenario is required to appear below an energy scale of $10^{10}$ GeV. Extending the validity of the model to higher energies requires heavier DM, with masses around $10^{10}$ GeV allowing the cutoff scale to reach values comparable to the GUT scale or a typical string scale. When a DM four-point self-coupling is included, achieving a cutoff scale at the GUT scale generally requires a heavier DM mass. In particular, when the self-coupling dominates the positivity bound, the DM mass is required to be approximately one order of magnitude larger, around $10^{11}$ GeV. We find that, in the presence of both interactions, achieving a cutoff scale at the GUT scale generally requires a DM mass of order $10^{10}$$10^{11}\,\mathrm{GeV}$ (or above), with the larger values favored when the four-point self-coupling plays a significant role.
In this heavy-mass regime of Higgs-portal DM the observed relic abundance can be explained by the freeze-in mechanism. The Higgs-portal coupling must be tiny, $\lambda_{h\phi}\lesssim 3.5\times 10^{-11}$. For DM masses of order $10^{10}$$10^{11}\,\mathrm{GeV}$ (or above), which are required to realize a GUT-scale cutoff, the reheating temperature is constrained to be $T_{\mathrm{reh}}\lesssim 10^{14}\,\mathrm{GeV}$.
@@ Acknowledgements
To obtain the relevant part of Eq. (7), we multiply $(s-2m^{2}_{\phi}+t/2)^{2}$ to the above Eq. (77). Here